# Prompt-based Personality Profiling: Reinforcement Learning for Relevance Filtering

### v1

@@ 1 Introduction
Our contributions are therefore:
@@ 2.1 Zero-Shot Predictions with Large Language Models
The terms prompt-based learning or in-context learning point at methods in which we use an LLM’s ability to generate text as a proxy for another task. This approach has proven effective for a variety of tasks (Yin et al., 2019; Gao et al., 2021; Cui et al., 2021; Ma et al., 2022; Sainz et al., 2021; Tu et al., 2022, i.a.). For example in a sentiment polarity classification, a classification instance could be combined with a prompt that requests a language model to output a word that corresponds either to a positive or a negative class (“The food is very tasty.” – “This review is positive/negative.”).
@@ 4.3 Baselines and Derived Systems
In this system, the selection process using the trained agent is replaced by selecting $N$ posts based on their relevance-score (Equation 5). Here, instances are ranked based on NPMI information and the top-N instances are directly provided to CNet. With this system, we aim to provide insights on performance of such a selection system when simply relying on information theoretic measures.
@@ 4.5 Results
Table 3: Macro $\textrm{F}_{1}$ (m-$\textrm{F}_{1}$) and weighted average $\textrm{F}_{1}$ (w-$\textrm{F}_{1}$) scores for all models on testing data (average of 10 runs with standard deviation). For models with top-$N$ parameter (lower part in this table), the best setting based on macro $\textrm{F}_{1}$ score on validation data is chosen for each trait (validation results shown in Table 5 in the Appendix).
@@
Table 3 shows the results. Here, we are in particular interested whether our approach is preferable compared to using all posts of profiles in a zero-shot setting. We therefore compare the third and fourth row in this Table and find that, for all traits except for extraversion, our approach (RL-Profiler) performs only slightly worse compared to using all posts (ALL+CNet). On average over all traits, we find that RL-Profiler performs worse by 1.8pp macro $\textrm{F}_{1}$ (46.3% vs. 48.1%) and 5.2pp weighted $\textrm{F}_{1}$ (78.9% vs. 73.7%). This is, although our method only uses 10 posts from each profile on average over all traits while the ALL+CNet system uses 92.9. We provide the number of selected posts for these methods for individual traits along with the corresponding prediction times in Appendix A.4.
@@
We find that our approach is preferable to selecting data at random when predicting personality, and only slightly worse compared to using all available posts of profiles. The advantage is that using only a small subset of posts increases efficiency of the zero-shot setting drastically. For example, when predicting extraversion, the average prediction time for a profile is reduced by more than 76% moving from 1.65s to 0.38s on the comparison between our method and the system using all available posts in a zero-shot setting. We provide further implementation details in the Appendix A.4.
@@ 4.6 Post-hoc Analysis with Artificial Data
Table 4: Macro $\textrm{F}_{1}$ (m-$\textrm{F}_{1}$) and weighted average $\textrm{F}_{1}$ (w-$\textrm{F}_{1}$) scores for all models on artificially enriched testing data (average of 10 runs with standard deviation). For models with top-$N$ parameter (lower part in this table), the best setting based on macro $\textrm{F}_{1}$ score on validation data is chosen for each trait (validation results shown in Table 6).
@@ 5 Conclusion and Future Work
We outlined a novel approach for automatic personality prediction from social media data which enables prompt-based predictions to focus on the most relevant parts of an input. Notably, we do not require labels of relevance, but induce the filter only from the prompt-performance on the profile level. While the results on real data shows no performance improvement overall, it does decrease the required context window of the language model. With experiment on artificial data, we can show a substantial performance improvement. This shows that our method helps the language model to focus on relevant content, instead of leaving this task to the attention mechanisms in the transformer.
@@ Acknowledgments
We acknowledge the support of the Ministerium für Wissenschaft, Forschung und Kunst Württemberg (MWK, Ministry of Science, Research and the Arts Baden-Württemberg under Az. 33-7533-9-19/54/5) in “Künstliche Intelligenz & Gesellschaft: Reflecting Intelligent Systems for Diversity, Demography and Democracy (IRIS3D)” and the support by the “Interchange Forum for Reflecting on Intelligent Systems” (IRIS) at the University of Stuttgart. Further, we acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). We further acknowledge support by the German Research Foundation (DFG) under project numbers KL 2869/1-2, KL 2869/11-1, KL 2869/12-1, and KL 2869/13-1.
@@ A.3 Artificial Post Generation and Dataset Enrichment
A.4 Implementation Details
We performed our experiments on a single NVIDIA RTX A6000 (48GB) GPU with AMD EPYC 7313 CPU.
We present the average prediction time per profile on testing data for different zero-shot systems in Table 9. For the RL-Profiler and RND+CNet system these times include time required to select a number of posts from profiles by using the trained agent or at random, respectively, and the time CNet requires to provide a prediction given the selected posts. For the ALL+CNet system this time only includes the time to retrieve a prediction from CNet. We find that runtime majorly depends on the number of posts selected by each system. Note that if selecting the same amount of posts from profiles, prediction time per profile for the PT+CNet and RL-Profiler systems are equal, since they use the same model architecture.
Variable RL-Profiler RND+CNet ALL+CNet
Openness 0.54 (5) 1.11 (50) 1.72 (94.3)
Conscientiousness 0.88 (5) 1.29 (50) 2.11 (93.7)
Extraversion 0.38 (5) 1.10 (50) 1.65 (92.4)
Agreeableness 0.61 (5) 1.12 (50) 1.57 (91.9)
Neuroticism 1.12 (30) 1.03 (30) 1.78 (92.1)
Table 9: Average prediction time in seconds per profile on testing data. For the RL-Profiler and RND+CNet system, the best setting for top-$N$ (in parentheses) based on validation performance is shown for each trait. Average number of posts if using all available posts from profiles in the given dataset.

### v2

@@ 1 Introduction
Our contributions are therefore [^2] The source code used in this study is available at: https://github.com/bluzukk/rl-profiler:
@@ 2.1 Zero-Shot Predictions with Large Language Models
The terms prompt-based learning or in-context learning point at methods in which we use an LLM’s ability to generate text as a proxy for another task. This approach has proven effective for a variety of tasks (Yin et al., 2019; Gao et al., 2021; Cui et al., 2021; Ma et al., 2022; Sainz et al., 2021; Tu et al., 2022, i.a.). For example, in a sentiment polarity classification, a classification instance could be combined with a prompt that requests a language model to output a word that corresponds either to a positive or a negative class (“The food is very tasty.” – “This review is positive/negative.”).
@@ 4.3 Baselines and Derived Systems
In this system, the selection process using the trained agent is replaced by selecting $N$ posts based on their relevance-score (Equation 5). Here, instances are ranked based on NPMI information and the top-$N$ instances are directly given to CNet. With this system, we aim to provide insights on performance of such a selection system when simply relying on information theoretic measures.
@@ 4.5 Results
Table 3: Macro $\textrm{F}_{1}$ (m-$\textrm{F}_{1}$) and weighted average $\textrm{F}_{1}$ (w-$\textrm{F}_{1}$) scores for all models on testing data (average of 10 runs with standard deviation). For models with top-$N$ parameter (lower part in this table), the best setting based on macro $\textrm{F}_{1}$ score on validation data is chosen for each trait (validation results with all settings for top-$N$ shown in Table 6).
@@
Table 3 shows the results. Here, we are in particular interested whether our approach is preferable compared to using all posts of profiles in a zero-shot setting. We therefore compare the third and fourth row in this Table and find that, for all traits except for extraversion, our approach (RL-Profiler) performs only slightly worse compared to using all posts (ALL+CNet). On average over all traits, we find that RL-Profiler performs worse by 1.8pp macro $\textrm{F}_{1}$ (46.3% vs. 48.1%) and 5.2pp weighted $\textrm{F}_{1}$ (78.9% vs. 73.7%). This is, although our method only uses 10 posts from each profile on average over all traits while the ALL+CNet system uses 92.9.
@@
Computational Analysis.
We perform our experiments on a single NVIDIA RTX A6000 (48GB) GPU with AMD EPYC 7313 CPU and present the average prediction time per profile on testing data for different zero-shot systems in Table 4. For the RL-Profiler and RND+CNet systems, the reported time includes both the time required to select a number of posts from each profile – using the trained agent or random selection, respectively – and the time taken by CNet to generate a prediction based on the selected posts. For the ALL+CNet system this time only reflects the duration required to retrieve a prediction from CNet. We find that prediction time is substantially reduced by a reduced number of selected posts. For example, when predicting extraversion, the average prediction time for a profile is reduced by more than 76% moving from 1.65s to 0.38s on the comparison between our method and the system using all available posts in a zero-shot setting.
Variable RL-Profiler RND+CNet ALL+CNet
Open. 0.54 (5) 1.11 (50) 1.72 (94.3)
Consc. 0.88 (5) 1.29 (50) 2.11 (93.7)
Extrav. 0.38 (5) 1.10 (50) 1.65 (92.4)
Agree. 0.61 (5) 1.12 (50) 1.57 (91.9)
Neur. 1.12 (30) 1.03 (30) 1.78 (92.1)
Table 4: Average prediction time in seconds per profile on testing data. For the RL-Profiler and RND+CNet system, the best setting for top-$N$ (in parentheses) based on validation performance is shown for each trait. For ALL+CNet the number in parentheses denotes the average number of posts available per profile.
@@
We find that our approach is preferable to selecting data at random when predicting personality, and only slightly worse compared to using all available posts of profiles. The advantage is that using only a small subset of posts increases efficiency of the zero-shot setting drastically.
@@ 4.6 Post-hoc Analysis with Artificial Data
Table 5: Macro $\textrm{F}_{1}$ (m-$\textrm{F}_{1}$) and weighted average $\textrm{F}_{1}$ (w-$\textrm{F}_{1}$) scores on artificially enriched testing data (average of 10 runs with standard deviation). For models with top-$N$ parameter (lower part), the best setting based on macro $\textrm{F}_{1}$ score on validation data is chosen for each trait (validation results with all settings for top-$N$ shown in Table 7).
@@ 5 Conclusion and Future Work
We outlined a novel approach for automatic personality prediction from social media data which enables prompt-based predictions to focus on the most relevant parts of an input. Notably, we do not require labels of relevance, but induce the filter only from the prompt-performance on the profile level. While the results on real data shows no performance improvement overall, it does decrease the required context window of the language model. With an experiment on artificial data, we can show a substantial performance improvement. This shows that our method helps the language model to focus on relevant content, instead of leaving this task to the attention mechanisms in the transformer.
@@ Acknowledgments
This paper is supported by the project INPROMPT (Interactive Prompt Optimization with the Human in the Loop for Natural Language Understanding Model Development and Intervention, funded by the German Research Foundation, KL 2869/13-1, project numer 521755488). We acknowledge the support of the Ministerium für Wissenschaft, Forschung und Kunst Württemberg (MWK, Ministry of Science, Research and the Arts Baden-Württemberg under Az. 33-7533-9-19/54/5) in “Künstliche Intelligenz & Gesellschaft: Reflecting Intelligent Systems for Diversity, Demography and Democracy (IRIS3D)” and the support by the “Interchange Forum for Reflecting on Intelligent Systems” (IRIS) at the University of Stuttgart. Further, we acknowledge the support by the Stuttgart Center for Simulation Science (SimTech).